OpenStudy (anonymous):

The length of a rectangle is 5x its width. If the area of a rectangle is 320 yd^2, what is its perimeter?

OpenStudy (anonymous):

this seems so easy... but im just spacing how to even make an equation out of this..

OpenStudy (anonymous):

320 = lw 320 (5w) w ?

OpenStudy (anonymous):

lets say the width is 'x'. Then the length is 5x since it is "5 times the width". The area of this rectangle would then be (Legth)(Width) = (5x)(x) :\[5x^{2}\] Setting this equal to the area of the rectangle (320 square yds) gives us: \[5x^{2}=320 \Rightarrow x^{2} = 64 \Rightarrow x = \pm 8\] but you cant have a "length" of negative 8, so x is 8. This means the width is 8 yrds, and the length is 5(8) = 40 yrds. The perimeter is 2(l + w) = 2(40+8) = 2(48) = 96 yrds

OpenStudy (anonymous):

ok! thanks!!!